# DynamicControl : Adaptive Condition Selection for Improved Text-to-Image Generation

### v2

@@ Abstract
Figure 1: (Left:) Multiple conditions generation results of our DynamicControl. (Right:) Comparison of different schemes dealing with multiple conditions in T2I task. (a) One condition is randomly selected, with an activated MOE encoder, (b) the input number of conditions is fixed manually, along with the pre-trained visual encoder, and (c) our proposed DynamicControl proposes a condition evaluator and multi-control adapter to select conditions adaptively.
@@ 1 Introduction
The emergence of generative diffusion models [8, 44, 14, 51, 52] has revolutionized image synthesis tasks, attributing to the significant enhancements in the quality and variety of generated images. Along with the large-scale image-text datasets [49, 48], the text-to-image (T2I) diffusion models have made further strides with Stable Diffusion Model (SDM) [44, 47]. Nonetheless, although linguistic control facilitates the creation of effective and engaging content, it still struggles to achieve fine-grained, pixel-level precision in controlling spatial, structural, and geometric aspects of the generated images. This presents significant challenges in the detailed manipulation of image attributes.
@@
Built upon ControlNet-like models [63, 27], various control signals such as layout constraints, segmentation maps, and depth maps have been explored to dictate the spatial arrangement, object shapes, and depth of field in generated images [35, 39, 61, 16, 65, 54]. Additionally, the field has witnessed the use of prompt engineering [29, 60, 64] and cross-attention constraints  [7, 21, 58] to further refine the regulation of image generation. Given the multiple conditions of a subject, one line (e.g. UniControl [39], Uni-ControlNet [65]) chooses to activate one condition at a time during the training process randomly, as shown in Fig. 1(a). This capacity to handle diverse visual conditions is quite inefficient and will greatly increase the computational burden and time costs of training. Another line of methods (e.g. AnyControl [54], ControlNet++ [27]) uses a fixed number (usually 2 or 4) of conditions and adopts MoE design or multi-control encoder to solve the varying-number conditions problem, as shown in Fig. 1(b). However, this fixed number scheme does not fundamentally solve the problem of multiple conditions, nor does it consider whether multiple conditions conflict with the generated results. While these methods have expanded the feasibility and applications of controlled image generation, a clear and comprehensive approach to enhance controllability under diverse conditions remains an area of ongoing research and development. This highlights the need for continued innovation in integrating and optimizing control mechanisms within T2I diffusion models to achieve more reliable and detailed image synthesis.
Given the multiple conditions from the same subject, as depicted in Fig. 2, it can be observed that for the same text prompt, different conditions generate different results in terms of color, texture, layout, rationality, etc. Moreover, from the SSIM score of the similarity with the source image, we can also see that different conditions struggle to accurately generate images that are consistent with the input source image. This also reveals that different conditions contribute differently to generating a better image, and some conditions even have a negative impact. Therefore, it is suboptimal to select only one or a fixed number of conditions in previous methods without considering their importance in generating an image closer to the source image and the internal relationship between each condition. To address this issue, we propose DynamicControl , a new framework that supports dynamic combinations of diverse control signals, which can adaptively select different numbers and types of conditions, as shown in Fig. 1(c).
Figure 2: Illustration of image generation results with different SSIM scores under different conditions.
@@
Specifically, we begin by designing a double-cycle controller that aims to generate the initial real score sorting for all the input conditions. Within the double-cycle controller, a pre-trained conditional generation model is utilized to generate an image based on each given image condition and text prompt, then we extract the corresponding image condition from the generated image using pre-trained discriminative models. Thus, the first cycle consistency is defined as the similarity between the extracted condition and each input condition. Furthermore, considering the pixel-level similarity of source image, the second cycle consistency is performed in the calculation of the similarity between the generated image and the source image. Combining the two similarity scores, this double-cycle controller will give the combined score ranking. However, this ranking requires generating initial images for all the conditions with random noise and the source image cannot be acquired during inference, which limits its full potential. To address these limitations, we introduce the Multimodal Large Language Model (MLLM) (e.g., LLaVA) [32, 69] into our model to build an efficient condition evaluator. This evaluator takes various conditions and promptable instructions as input and optimizes the best ordering of the conditions with the score ranking from the double-cycle controller. With a dynamic selection scheme, the final sorting results from the pre-trained condition evaluator are fed into the parallel multi-control adapter to learn necessary different level feature maps from dynamic visual conditions, where unique information from different visual conditions is captured adaptively. In this way, only those control conditions that are harmonious and mutually advantageous to the generated results are preserved. The output embeddings can be integrated to modulate ControlNet [63], facilitating task-specific visual conditioning controls. Consequently, our DynamicControl promotes enhanced and more harmonious control over the generated images.
Our main contributions are summarized as follows:
@@ 2 Related Work
Text-to-image Generation. Text-to-Image (T2I) diffusion models [36, 43, 44, 47] have rapidly evolved as a leading approach for generating high-quality images from textual prompts, offering a fresh perspective on image synthesis [23, 13, 8]. Initially rooted in image generation, diffusion models [14, 51] have been adeptly tailored to the T2I domain, utilizing a process that incrementally introduces and then removes noise, allowing for the progressive refinement of image quality [38, 42, 45, 41]. This iterative denoising process, coupled with the ability to condition on both text inputs and intermediate image representations, enhances control over the generation process. Recent advancements [9, 34, 3, 22, 5, 17] in T2I diffusion models have incorporated various techniques to improve alignment between textual and visual features. Other models, including notable variants like DALLE-2 [43] and Stable Diffusion [44], have demonstrated superior capability in capturing fine-grained structures and textures compared to earlier generative approaches. Stable Diffusion, in particular, has scaled up the latent diffusion approach with larger models and datasets, making these models accessible to the public.
Controllable Image Synthesis. To achieve fine-grained control over generated images, text descriptions alone often fall short in providing detailed guidance, necessitating the integration of diverse modalities for enhanced control. For instance, instance-based controllable generation methods [57, 68] allow for location control through more free-form inputs like points, scribbles, and boxes, while structure signals like sketches and depth maps further refine the visual output.
Recent advancements have seen the development of frameworks like ControlNet [63], ControlNet++ [27] and T2I-Adapter [35], which incorporate trainable modules within T2I diffusion models to encode additional control signals into latent representations. Moreover, unified models [16, 39, 46, 54] akin to ControlNet have been proposed to handle multiple control signals within a single framework, supporting multi-control image synthesis. These models typically use fixed-length input channels or a mixture of experts (MoE) design with hand-crafted weighted summation to aggregate conditions effectively. However, despite these innovations, challenges persist in managing conditions with complex interrelations and achieving harmonious, natural results under varied control signals.
@@
MLLM with Diffusion Models. Recent advancements in Vision Large Language Models (VLLMs) have significantly enhanced the performance of vision tasks, leveraging the extensive world knowledge and complex instruction comprehension capabilities of these models [2, 30, 32, 31, 6]. Notably, the open-sourced LLaMA model [55] has been instrumental in improving image-text alignment through instruction-tuning, a technique further refined by models such as LLaVA [31] and MiniGPT-4 [32, 69]. These models have demonstrated robust capabilities across a variety of tasks, particularly those reliant on text generation. In the realm of image generation, fine-tuning VLLMs has shown great success [12, 28, 25, 10, 11]. For instance, SmartEdit [19] adapts the LLaVA model to specialize in image editing tasks. FlexEdit [56] employs a VLLM in comprehending the image content, mask, and user instructions. Additionally, models like Emu [53] and CM3Leon [62] have expanded the capabilities of multi-modal language models, employing architectures and training methods adapted from text-only models to execute both text-to-image and image-to-text generation tasks effectively.
@@ 3 Method
In this section, we first introduce the background of diffusion models in Section 3.1. The pipeline of DynamicControl is demonstrated in Fig. 3. Given the multiple conditions, we first introduce the double-cycle controller (Section 3.2) to produce the real ranking score as the supervision signal for training the condition evaluator (Section 3.3) combined with MLLMs. Then, these ranked conditions with selection scores from the pre-trained condition evaluator are dynamically encoded by the multi-control adapter (Section 3.4) to fulfill controllable image generation. Finally, we discuss how to jointly optimize MLLMs with diffusion models and train our multiple conditional T2I model (Section 3.5).
3.1 Preliminary
Diffusion models, as delineated in [14], establish a Markovian sequence known as the forward diffusion process $q(x_{t}|x_{0})$, which incrementally introduces noise into the initial data $x_{0}$:
$x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,\quad\epsilon\sim\mathcal{N}(\mathbf{0},{I}),$ (1)
where $\epsilon$ denotes a noise vector drawn from a Gaussian distribution, and $\bar{\alpha}_{t}$ is defined as $\bar{\alpha}_{t}:=\prod_{s=0}^{t}\alpha_{s}$. Here, $\alpha_{t}=1-\beta_{t}$ represents a function of the timestep $t$, which is specified by the denoising algorithm such as DDPM [14]. Consequently, the diffusion training loss is expressed as:
$\footnotesize\mathcal{L}\left(\epsilon_{\theta}\right)=\sum_{t=1}^{T}\mathbb{E}_{x_{0}\sim q\left(x_{0}\right),\epsilon\sim\mathcal{N}(\mathbf{0},{I})}\left[\left\|\epsilon_{\theta}\left(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon\right)-\epsilon\right\|_{2}^{2}\right].$ (2)
In scenarios of controllable generation, where specific image conditions $c_{v}$ and text prompts $c_{t}$ are provided, the diffusion training loss at timestep $t$ is reformulated as:
$\small\mathcal{L}_{\text{diffusion}}=\mathbb{E}_{x_{0},t,c_{t},c_{v},\epsilon\sim\mathcal{N}(0,1)}\left[\|\epsilon_{\theta}\left(x_{t},t,c_{t},c_{v}\right)-\epsilon\|_{2}^{2}\right].$ (3)
During the inference, starting from a randomly sampled noise $x_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, the final denoised image $x_{0}$ is predicted through a sequential denoising process:
$x_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(x_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}{\epsilon}_{\theta}\left(\mathbf{x}_{t},t\right)\right)+\sigma_{t}\epsilon,$ (4)
where $\epsilon_{\theta}$ signifies the noise predicted at timestep $t$ by UNet [45] model parameterized by $\theta$, and $\sigma_{t}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}$ is the variance of posterior Gaussian distribution $p_{\theta}(x_{0})$.
@@ 3.2 Double-Cycle Controller
To be more specific, given the multiple conditions along with text prompts, we first utilize a pre-trained conditional generation model [63, 27] to generate images for each condition. Then corresponding reverse conditions are extracted by different pre-trained discriminative models. Based on these generated images and reverse conditions, we design a double-cycle controller to make an initial importance assessment of the input multiple control conditions. This double-cycle controller consists of two consistency scores, namely condition consistency and image consistency.
Condition Consistency. Inspired by  [70, 27], for each input condition $c_{i,v}$ ($i=1,2,...,N$, $N$ is the total number of conditions) and the corresponding output condition $\hat{c}_{i,v}$ of the generated image $x_{0}^{\prime}$, we optimize the condition cycle consistency loss for better controllability, which is formulated as:
@@
Here we perform single-step sampling [14] on disturbed image $x^{\prime}_{t}$, which means $x_{0}\approx x^{\prime}_{0}=\frac{x^{\prime}_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}\left(x^{\prime}_{t},c_{i,v},c_{i,t},t\right)}{\sqrt{\alpha_{t}}}$, where $\mathbb{D}$ is the discriminative reward model to optimize the controllability of $\mathbb{G}$. $\mathcal{L}$ represents an abstract metric function that is adaptable to various concrete forms depending on specific visual conditions. This flexibility allows it to be tailored to meet the unique requirements of different visual analysis tasks, enhancing the applicability and effectiveness of the model across diverse scenarios.
Reverse Image Consistency. Apart from the condition consistency, we employ a reverse image consistency loss to guarantee that the original image is similar to the generated one. We achieve this by minimizing pixel-wise and semantic discrepancies between the generated image and the source image. Given the CLIP embeddings [40] of the source image $E_{I_{source}}$, generated image $E_{I_{gen}}$, the loss is defined as:
@@ 3.3 Condition Evaluator
As shown in Fig. 3, given the conditions $c_{1},c_{2},\ldots,c_{N}$ and instruction $\tau$, our primary objective is to optimize the best ordering of the conditions with the score ranking from the double-cycle controller. Inspired by  [24, 18], we expand the original LLM vocabulary of LLaVA [32] with $N$ new tokens “<$con^{0}$>, …, <$con^{N}$>” to represent generation information and append these tokens to the end of instruction $\tau$. Then, the conditions $c_{1},c_{2},\ldots,c_{N}$ and the reorganized instruction $\tau^{\prime}$ are fed into the Vision Large Language Model (VLLM) $LLaVA(\cdot;\omega)$ to obtain response tokens, which are processed to extract the corresponding hidden states $h_{i}\in\mathcal{H}$, capturing the deeper semantic information from the VLLM’s representations of the inputs. However, these hidden states predominantly exist within the text vector space of the LLM, presenting compatibility issues when interfacing with a diffusion model, especially one trained on CLIP text embeddings [40]. This discrepancy can hinder effective integration between the models. Considering this, we transfer Q-Former [26] to refine the hidden states into embeddings $f_{c}$ compatible with the diffusion model. The transformation process is represented as:
$R{}$ $=LLaVA(c_{1},c_{2},...,c_{N},\tau^{\prime};\omega),$ (7)
$h_{i}$ $=H(c_{1},c_{2},...,c_{N},\tau^{\prime};\omega|r_{i}),$
@@
[…]
@@
Subsequently, the predicted results from the LLM for each condition are supervised by corresponding ranking scores from double-cycle controller, optimizing the final sorting rankings. The process is represented as:
$\mathcal{L}_{\text{eval}}=-\sum_{i=1}^{N}c_{i}\log p_{i}.$ (9)

### v3

@@ Abstract
Figure 1: Visualization comparison between method of fixed number control conditions (Uni-ControlNet (Zhao et al., 2024)) and our proposed method in different conditional controls with the same text prompt. (a, left two columns) Text and various visual controls, where $C_{1}$, $C_{2}$, $C_{3}$ and $C_{4}$ denotes different control conditions. (b, middle three or four columns) Generation results from UniControlNet. (c, last column) Generation results from our DynamicControl. Previous methods struggled to generate coherent results under multiple conditions, while our results maintain strong similarity to the respective visual controls.
@@ 1 Introduction
The emergence of generative diffusion models (Dhariwal & Nichol, 2021; Rombach et al., 2022; Ho et al., 2020; Song et al., 2020a; b) has revolutionized image synthesis tasks, attributing to the significant enhancements in the quality and variety of generated images. Built upon ControlNet-like models (Zhang et al., 2023; Li et al., 2025), various control signals such as layout constraints, segmentation maps, and depth maps have been explored to dictate the spatial arrangement, object shapes, and depth of field in generated images (Mou et al., 2024; Qin et al., 2023; Ye et al., 2023; Hu et al., 2023; Zhao et al., 2024; Sun et al., 2024). As noted in (Sun et al., 2024), different visual control signals have complementary properties. For instance, depth maps can effectively govern spatial relationships between objects but fall short in capturing fine-grained object details, while canny maps excel at capturing precise texture contours yet overlook the global structural context. In practical use cases, it is often necessary to describe the visual features of a key object through multiple visual conditions to achieve accurate control over its generation. Users generally aim to control both the overall layout and intricate details simultaneously. However, integrating visual conditions that contain rich layout and detail information into a single visual condition map remains a challenge.
Given the multiple conditions of a subject, one line (e.g. UniControl (Qin et al., 2023), Uni-ControlNet (Zhao et al., 2024)) chooses to activate one condition at a time during the training process randomly. This capacity to handle diverse visual conditions is quite inefficient and will greatly increase the computational burden and time costs of training. Another line of methods (e.g. AnyControl (Sun et al., 2024), ControlNet++ (Li et al., 2025)) uses a fixed number (usually 2 or 4) of conditions and adopts MoE design or multi-control encoder to solve the varying-number conditions problem.
@@
However, this fixed number scheme does not fundamentally solve the problem of multiple conditions, nor does it consider whether multiple conditions conflict with the generated results. As shown in Fig. 1(b), it is suboptimal to select only one or a fixed number of conditions in previous methods without considering their importance in generating an image closer to the source image and the internal relationship between each condition. While these methods have expanded the feasibility and applications of controlled image generation, a clear and comprehensive approach to enhance controllability under diverse conditions remains an area of ongoing research and development. This highlights the need for continued innovation in integrating and optimizing control mechanisms within T2I diffusion models to achieve more reliable and detailed image synthesis.
To address this issue, we propose DynamicControl , a new framework that supports dynamic combinations of diverse control signals, which can adaptively select different numbers and types of conditions, facilitating more harmonious and natural generation results, as shown in Fig. 1(c).
@@
Specifically, we begin by designing a double-cycle controller that aims to generate the initial real score sorting for all the input conditions. Within the double-cycle controller, a pre-trained conditional generation model is utilized to generate an image based on each given image condition and text prompt, then we extract the corresponding image condition from the generated image using pre-trained discriminative models. Thus, the first cycle consistency is defined as the similarity between the extracted condition and each input condition. Furthermore, considering the pixel-level similarity of source image, the second cycle consistency is performed in the calculation of the similarity between the generated image and the source image. Combining the two similarity scores, this double-cycle controller will give the combined score ranking. However, this ranking requires generating initial images for all the conditions with random noise and the source image cannot be acquired during inference, which limits its full potential. To address these limitations, we introduce the Multimodal Large Language Model (MLLM) (e.g., LLaVA) (Liu et al., 2024b; Zhu et al., 2023) into our model to build an efficient condition evaluator. This evaluator takes various conditions and promptable instructions as input and optimizes the best ordering of the conditions with the score ranking from the double-cycle controller. With a dynamic selection scheme, the final sorting results from the pre-trained condition evaluator are fed into the parallel multi-control adapter to learn necessary different level feature maps from dynamic visual conditions, where unique information from different visual conditions is captured adaptively. In this way, only those control conditions that are harmonious and mutually advantageous to the generated results are preserved. The output embeddings can be integrated to modulate ControlNet (Zhang et al., 2023), facilitating task-specific visual conditioning controls. Consequently, our DynamicControl promotes enhanced and more harmonious control over the generated images. Our main contributions are summarized as follows:
@@ 2 Related Work
Text-to-image Generation. Text-to-Image (T2I) diffusion models (Nichol et al., 2021; Ramesh et al., 2022; Rombach et al., 2022; Saharia et al., 2022) have rapidly evolved as a leading approach for generating high-quality images from textual prompts, offering a fresh perspective on image synthesis (Kingma et al., 2021; Ho & Salimans, 2022; Dhariwal & Nichol, 2021). Initially rooted in image generation, diffusion models (Ho et al., 2020; Song et al., 2020a) have been adeptly tailored to the T2I domain, utilizing a process that incrementally introduces and then removes noise, allowing for the progressive refinement of image quality (Podell et al., 2023; Ramesh et al., 2021; Ronneberger et al., 2015; Raffel et al., 2020). This iterative denoising process, coupled with the ability to condition on both text inputs and intermediate image representations, enhances control over the generation process. Recent advancements (Gal et al., 2022; Meng et al., 2021; Brooks et al., 2023; Kawar et al., 2023; Cao et al., 2023; Huang et al., ) in T2I diffusion models have incorporated various techniques to improve alignment between textual and visual features. Other models, including notable variants like DALLE-2 (Ramesh et al., 2022) and Stable Diffusion (Rombach et al., 2022), have demonstrated superior capability in capturing fine-grained structures and textures compared to earlier generative approaches. Stable Diffusion, in particular, has scaled up the latent diffusion approach with larger models and datasets, making these models accessible to the public.
Controllable Image Synthesis. To achieve fine-grained control over generated images, text descriptions alone often fall short in providing detailed guidance, necessitating the integration of diverse modalities for enhanced control. For instance, instance-based controllable generation methods (Wang et al., 2024b; Zhou et al., 2024) allow for location control through more free-form inputs like points, scribbles, and boxes, while structure signals like sketches and depth maps further refine the visual output.
Recent advancements have seen the development of frameworks like ControlNet (Zhang et al., 2023), ControlNet++ (Li et al., 2025) and T2I-Adapter (Mou et al., 2024), which incorporate trainable modules within T2I diffusion models to encode additional control signals into latent representations. Moreover, unified models (Huang et al., 2022; Ham et al., 2023; Hu et al., 2023; Qin et al., 2023; Sun et al., 2024) akin to ControlNet have been proposed to handle multiple control signals within a single framework, supporting multi-control image synthesis. These models typically use fixed-length input channels or a mixture of experts (MoE) design with hand-crafted weighted summation to aggregate conditions effectively. However, despite these innovations, challenges persist in managing conditions with complex interrelations and achieving harmonious, natural results under varied control signals. Bridging the redundant information among these visual conditions and reasonably utilizing the emphasis of various visual conditions to coordinate the generation of the object is exactly the issue that this paper wants to focus on.
@@
MLLM with Diffusion Models. Recent advancements in Vision Large Language Models (VLLMs) have significantly enhanced the performance of vision tasks, leveraging the extensive world knowledge and complex instruction comprehension capabilities of these models (Bai et al., 2023; Lin et al., 2023; Liu et al., 2024b; a; Chen et al., 2023). Notably, the open-sourced LLaMA model (Touvron et al., 2023) has been instrumental in improving image-text alignment through instruction-tuning, a technique further refined by models such as LLaVA (Liu et al., 2024a) and MiniGPT-4 Liu et al. (2024b); Zhu et al. (2023). These models have demonstrated robust capabilities across a variety of tasks, particularly those reliant on text generation. In the realm of image generation, fine-tuning VLLMs has shown great success (Ge et al., 2024; Li et al., 2023b; Koh et al., 2024; Ge et al., 2023a; b). For instance, SmartEdit (Huang et al., 2024b) adapts the LLaVA model to specialize in image editing tasks. FlexEdit (Wang et al., 2024a) employs a VLLM in comprehending the image content, mask, and user instructions. Additionally, models like Emu (Sun et al., 2023) and CM3Leon (Yu et al., 2023) have expanded the capabilities of multi-modal language models, employing architectures and training methods adapted from text-only models to execute both text-to-image and image-to-text generation tasks effectively.
@@ 3 Method
The pipeline of DynamicControl is demonstrated in Fig. 2. Given the multiple conditions, we first introduce the double-cycle controller (Section 3.1) to produce the real ranking score as the supervision signal for training the condition evaluator (Section 3.2) combined with MLLMs. Then, these ranked conditions with selection scores from the pre-trained condition evaluator are dynamically encoded by the multi-control adapter (Section 3.3) to fulfill controllable image generation. Finally, we discuss how to jointly optimize MLLMs with diffusion models and train our multiple conditional T2I model (Section 3.4).
@@ 3.1 Double-Cycle Controller
To be more specific, given the multiple conditions along with text prompts, we first utilize a pre-trained conditional generation model (Zhang et al., 2023; Li et al., 2025) to generate images for each condition. Then corresponding reverse conditions are extracted by different pre-trained discriminative models. Based on these generated images and reverse conditions, we design a double-cycle controller to make an initial importance assessment of the input multiple control conditions. This double-cycle controller consists of two consistency scores, namely condition consistency and image consistency.
Condition Consistency. Inspired by  (Zhu et al., 2017; Li et al., 2025), for each input condition $c_{i,v}$ (i=1,2,,Ni=1,2,...,N, NN is the total number of conditions) and the corresponding output condition $\hat{c}_{i,v}$ of the generated image $x_{0}^{\prime}$, we optimize the condition cycle consistency loss for better controllability, which is formulated as:
@@
Here we perform single-step sampling (Ho et al., 2020) on disturbed image $x^{\prime}_{t}$, which means $x_{0}\approx x^{\prime}_{0}=\frac{x^{\prime}_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}\left(x^{\prime}_{t},c_{i,v},c_{i,t},t\right)}{\sqrt{\alpha_{t}}}$, where $\mathbb{D}$ is the discriminative reward model to optimize the controllability of $\mathbb{G}$. $\mathcal{L}$ represents an abstract metric function that is adaptable to various concrete forms depending on specific visual conditions. This flexibility allows it to be tailored to meet the unique requirements of different visual analysis tasks, enhancing the applicability and effectiveness of the model across diverse scenarios.
Reverse Image Consistency. Apart from the condition consistency, we employ a reverse image consistency loss to guarantee that the original image is similar to the generated one. We achieve this by minimizing pixel-wise and semantic discrepancies between the generated image and the source image. Given the CLIP embeddings (Radford et al., 2021) of the source image $E_{I_{source}}$, generated image $E_{I_{gen}}$, the loss is defined as:
@@ 3.2 Condition Evaluator
As shown in Fig. 2, given the conditions $c_{1},c_{2},\ldots,c_{N}$ and instruction $\tau$, our primary objective is to optimize the best ordering of the conditions with the score ranking from the double-cycle controller. Inspired by  (Koh et al., 2023; Huang et al., 2024a), we expand the original LLM vocabulary of LLaVA (Liu et al., 2024b) with $N$ new tokens $con^{0}$>, …, $con^{N}$>” to represent generation information and append these tokens to the end of instruction $\tau$. Then, the conditions $c_{1},c_{2},\ldots,c_{N}$ and the reorganized instruction $\tau^{\prime}$ are fed into the Vision Large Language Model (VLLM) $LLaVA(\cdot;\omega)$ to obtain response tokens, which are processed to extract the corresponding hidden states $h_{i}\in\mathcal{H}$, capturing the deeper semantic information from the VLLM’s representations of the inputs. However, these hidden states predominantly exist within the text vector space of the LLM, presenting compatibility issues when interfacing with a diffusion model, especially one trained on CLIP text embeddings (Radford et al., 2021). This discrepancy can hinder effective integration between the models. Considering this, we transfer Q-Former (Li et al., 2023a) to refine the hidden states into embeddings $f_{c}$ compatible with the diffusion model. The transformation process is represented as:
$R$ $=LLaVA(c_{1},c_{2},.,c_{N},\tau^{\prime};\omega),$ (3)
$h_{i}$ $=H(c_{1},c_{2},.,c_{N},\tau^{\prime};\omega|r_{i}),$
@@
where $r=\{\text{``\textless$con^{0}$\textgreater, \ldots, \textless$con^{N}$\textgreater''}\}\in R$ is the condition tokens set and $f_{c}$ represents the transformed embeddings by the Q-Former function $Q$. For fine-tuning efficiency, we utilize the LoRA (Hu et al., 2021) scheme, where the majority of parameters $\theta$ in the LLM are kept frozen. The recurrent optimization process can be formulated as:
@@
Subsequently, the predicted results from the LLM for each condition are supervised by corresponding ranking scores from double-cycle controller, optimizing the final sorting rankings. The process is represented as $\mathcal{L}_{\text{eval}}=-\sum_{i=1}^{N}c_{i}\log p_{i}$.